# Layered, Overlapping, and Inconsistent: A Large-Scale Analysis of the Multiple Privacy Policies and Controls of U.S. Banks

### v1

@@ 5.2. RQ2: Third-Party Sharing Statements
To show inconsistencies in marketing-related sharing statements across policies, we first discuss the GLBA statements (5.2.1). After presenting relevant statement categories we identified in other policies (5.2.2), we compare banks’ denying GLBA statements (5.2.3) and affirmative GLBA statements (5.2.4) each with opposite statements in other polices. Last, we point out a common misleading language use for no-sharing (5.2.5).
@@ 5.2.1. GLBA Third-Party Sharing Statements
Of the seven sharing purposes prescribed in the GLBA model notice’s data-sharing table (see Figure 1(a)), we focused on the four relevant to third-party sharing: 1,542 banks (77% out of 2002 banks that provided this row) stated they are sharing consumer personal information for marketing (marketing), 879 (43.9% out of 2002) disclosed sharing with nonaffiliated financial companies for joint marketing (joint marketing), 329 (16.4% out of 1070) disclosed sharing with affiliates for affiliates’ marketing (affiliate sharing), and 65 (3.2% out of 1925) disclosed sharing with nonaffiliates for nonaffiliates’ marketing (nonaffiliate sharing). A few banks (36, 1.8% out of 2002 banks with a GLBA notice) shared data for all four marketing purposes, 10.6% (214) for three of them; a third for two purposes (643, 32.1%) and another third shared for one of these purposes (743, 37.1%). Only 5.6% (112) indicated not sharing for any of the four purposes. These four GLBA-defined third-party sharing practices have nuanced differences, however, for consumers, they all describe scenarios in which their personal financial information is shared by their bank with other entities for marketing use.
Some banks include additional data-sharing disclosures in their GLBA notices’ OII box. Among the 718 banks that had a non-empty OII box, 332 included California-specific language. 300 of them provided statements further limiting data-sharing practices for California residents. The most common phrasing was: “We will not share personal information with non-affiliates either for them to market to you or for joint marketing without your authorization. We will also limit our sharing of personal information about you with our affiliates to comply with all California privacy laws that apply to us.”—which suggests that without consent, the bank dooes not share for joint marketing or nonaffiliate marketing for Californians, though remaining vague on affiliate sharing. 20 (of 332) banks included more minimal or opaque disclosures that mentioned CalFIPA but only stated “In accordance with California law, [bank] does not share personal information we collect except as permitted by law.” 55 (of 332) mentioned or linked to the bank’s CCPA policy (e.g., “Please visit our California Privacy Rights Act Policy for more information.”). These references at least acknowledge the existence of additional privacy policies, but still requires consumers to navigate beyond the standardized data-sharing table in the GLBA notice, into a less prominent section of the notice, follow a link, and then read those additional policies to fully understand the bank’s data-sharing practices.
@@ 5.2.2. Third-Party Sharing Statements in Other Privacy Policy Types
Thus, for the purpose of comparing across policies, we created statement categories based on whether a third-party sharing statement refers to personal information in general or to a specific data type, and whether it indicates sharing with external entities for marketing-related purposes. These categories include both statements allowing/describing third-party sharing (“yes”) and explicitly denying sharing (“no”). However, in contrast to GLBA disclosures that require explicit “no” statements, most banks’ policies are silent on sharing practices the bank presumably does not engage in. Appendix LABEL:appendix:disclosure-categories provides examples for each of the sharing statement types discussed below. We treat marketing and advertising as similar purposes, hereafter referred to as marketing.
@@
We assessed when banks indicated a “no” for each of four GLBA marketing-related sharing purposes (marketing, joint marketing, affiliate sharing, nonaffiliate sharing), how many banks simultaneously disclosed a “yes” to third-party sharing in related categories identified above. We found that many banks still disclosed varying amounts of sharing in other policies (see Table 2).
@@
Inconsistencies regarding nonaffiliate sharing. 1860 banks (92.9% of 2002 banks with an analyzable GLBA notice) stated in their GLBA notice that they do not share with nonaffiliates. Yet 494 of them (26.6%) also had affirmative sharing statements in other policies under at least one category we identified. Among these 494 banks, about 30% stated sharing personal information for advertising purposes, more than two-thirds disclosed using third-party analytics cookies, and about 40% for third-party marketing cookies. About 20% disclosed sharing personal information for marketing in CCPA-related disclosures. Possibly less concerning but still an indication of nonaffiliate sharing, less than 20% stated sharing for analytics purposes. More than 10% disclosed sharing specific personal information types for marketing (e.g., “[w]e may also share your device’s physical location, combined with information about what advertisements you viewed and other information we collect, with our marketing partners […]”). A small percentage, but still 17 banks in number, even disclosed selling data under CCPA (e.g., “In the preceding twelve (12) months, we have sold personal information”).
@@
These inconsistencies may reflect GLBA’s narrow scope of personal information as relating to financial information only, whereas privacy policies often cover a broader range of practices, including data collected through a bank’s website and mobile apps. Under GLBA, a “no” to nonaffiliate sharing means that the bank does not share (financial) personal information with nonaffiliates for those nonaffiliates’ marketing purposes. However, this does not preclude sharing with nonaffiliates for other purposes like supporting the bank’s own marketing efforts. This discrepancy is concerning as it may mislead consumers. A consumer may conclude from the GLBA notice that a bank does not share their personal information with others, while in practice (with a 26.5% likelihood) the bank still shares lots of their personal information with third parties.
@@
Inconsistencies regarding affiliate sharing. A similar definitional gap exists for affiliate sharing, which under GLBA is confined to sharing personal information with affiliated companies (e.g., a bank’s affiliated investment or loans business) for their marketing. Among the 741 banks that stated they do not share for affiliate marketing in their GLBA notice, 23.9% of them (177 banks) still disclosed other data-sharing practices in other policies. More than two-thirds of these 177 banks disclosed that they allow analytics cookies, more than a third disclosed marketing cookies, and about a quarter stated sharing for marketing.
@@
Inconsistencies regarding other sharing. We have seen above that when inconsistencies occur, most of them are between GLBA and third-party cookie disclosures. This pattern remains the same for “no” responses to GLBA-defined joint marketing and general marketing. Among the 1123 banks that said they do not share for joint marketing in GLBA notice, 263 banks (23.4% of them) in fact stated that they do share personal information for marketing purposes in other privacy policies. For those with at least one inconsistency, about 70% banks stated allowing third-party analytics cookies, and a third for marketing cookies. About a quarter of them stated sharing personal information for marketing. A smaller group of 460 banks indicated “no” to sharing for GLBA-defined marketing purpose, and 68 (14.8% of them) disclosed at least one sharing practice in other policies. Among the 68 banks, about 85% disclosed allowing third-party advertising or marketing cookies. Compared to other purposes, under joint marketing and marketing, we found smaller percentages of banks that had at least one inconsistency. Yet, any inconsistencies under the two purposes are striking, since they are about sharing for the bank’s own marketing purposes, a practice likely more common than nonaffiliate or affiliate sharing that is more narrowly defined under GLBA.
@@
Taken together, of the banks that indicated “no” to at least one GLBA-defined sharing, 505 also disclosed at least one data sharing practice as occurring in other policies. They account for 55.2% of the 915 banks that provided a GLBA and at least one additional policy. This suggests that even when banks claim not to share under GLBA categories, they are very likely to still engage in similar data-sharing practices for marketing. In sum, inconsistency between GLBA’s and other policies’ disclosures is widespread, which highlights coverage gaps and the lack of compatible clarity across privacy disclosures.
@@ 5.2.4. More Restrictive Third-Party Sharing Than Affirmative GLBA Statements
Mkt 37 43 117 188 85 329
Joint-mkt 19 32 65 99 50 193
Aff-mkt 7 16 21 87 36 120
Nonaff-mkt 1 5 6 14 4 20
Possibly less concerning, some banks that indicated they share personal information under GLBA-defined categories simultaneously stated in their other privacy policies that they do not share information with third parties—most commonly in their CCPA-related sharing disclosures (see Table 3).
Among the 329 banks that disclosed affiliate sharing (16.4% of 2002 banks with an analyzable GLBA notice), 120 banks (36.5%) had at least one statement denying data-sharing in their other policies. 72.5% percent of the 329 banks (87) stated that they don’t share personal information, with or without reference to CCPA definition of sharing in CCPA policies or California-specific sections, and thirty percent stated that they don’t sell personal information in CCPA-related disclosures. About twenty percent disclosed allowing marketing or analytics third-party cookies.
Regarding the GLBA’s nonaffiliate sharing, similarly high percentages of banks disclosed at least one denying data-sharing statement in their other policies, though fewer banks in number did so. Among the 65 banks (3.2% of 2002 banks with an analyzable GLBA notice) that stated a “no” to nonaffiliate sharing, 30.1% of them (20 banks) disclosed a denial of data-sharing of some kind. Seventy percent of the 20 banks indicated that they don’t share, and twenty percent don’t sell, in CCPA-related disclosures. Thirty percent stated allowing third-party marketing or analytics cookies.
Among banks that indicated a “yes” to GLBA’s marketing or joint-marketing purposes, smaller percentages (about 20% for each), though with larger counts, disclosed a denying data-sharing practice elsewhere. For both, more than half of the ones with inconsistency disclosed not sharing or selling in CCPA-related disclosures.
This pattern of discrepancy between GLBA and CCPA-related disclosures may indicate that banks specifically restrict their third-party sharing practices for their Californian customers while sharing their other customers’ data more freely.
@@ 5.2.5. Misleading “We Don’t Share” Statements
Figure 4. Data-sharing statements and indication of whether consumers can limit the corresponding sharing practice in GLBA notices (n=2002).
RQ2 Summary. Banks’ affirmative sharing disclosures under the four GLBA marketing-related categories ranged widely, from 77% to just 3%. Comparing these with disclosures in other privacy policies reveals persistent, concerning inconsistencies: 15% to 27% of banks said “no” under GLBA while saying “yes” elsewhere for related sharing. In total, 505 banks (55.2% of those with both a GLBA and other policies) show such inconsistencies. This highlights that GLBA “no” statements alone are insufficient for understanding banks’ data practices, especially when online and mobile services are involved. Additionally, some banks that disclosed sharing under GLBA restrict such practices specifically for California residents.
@@ 5.3. RQ3: Third-Party Sharing Opt-Outs
Table 4. Summary of opt-out controls provided (n=2073).
@@ 4 Other combinations 4 4
Under the GLBA, banks that share for nonaffiliate marketing, affiliate marketing, or affiliate credit-related purposes must provide respective opt-outs. We found that almost all banks followed these requirements and provided required GLBA opt-outs (see Figure 4). For sharing purposes without required opt-outs, few banks that disclosed sharing provided opt-outs (joint marketing (6.0%), affiliate transactions (4.5%), bank’s own marketing (2.9%) and everyday business (0.1%)). This indicates that banks generally only offer opt-out controls when mandated by law. There is also a risk that banks may categorize their data sharing practices as “joint marketing” rather than sharing for affiliate or nonaffiliate marketing to sidestep GLBA opt-out obligations.
Furthermore, we found that banks’ GLBA-related opt-out options are often burdensome for consumers to exercise. Calling the bank was the most commonly mentioned opt-out method (307 of 367 banks providing GLBA-related opt-outs, 83.7%), and for many the only method provided (114, 31.1% of 367). Additionally, 43 banks (11.7% of 367) required consumers to submit opt-out requests by mail, with no other methods offered. Only a third of banks (139, 37.9% of 367) offered a link to their webpage. Thus, while most banks technically comply with GLBA opt-out requirements, the opt-out methods are often unnecessarily cumbersome, potentially discouraging consumers from exercising their privacy rights.
@@ 5.3.3. Third-Party Cookie Controls
300 banks’ privacy policies mentioned that consumers can opt-out their use of cookies, but through browsers, e.g., “You can set your browser to refuse Cookies.” 106 banks in policies referred consumers to cookie opt-outs offered by third parties (e.g., the Network Advertising Initiative), often for advertising cookies specifically. Only 17 banks mentioned using their cookie setting/banner as an opt-out choice in policies.
We examined cookie controls implemented/provided on websites. Among the 332 banks that indicated allowing third-party cookies for marketing, analytics, or both purposes in policies, only 43 (13.0%) of them implemented cookie control (see Table 6). On the other hand, among the vast majority of banks (1,741) that did not disclose cookie practices in privacy policies, 53 (3.0% of 1,741) still implemented cookie controls. This suggests that they are engaging in cookie-related data sharing, which potentially involves third-party sharing, despite not explicitly disclosing it. Our analysis further revealed substantial variance in how cookie controls are presented on bank websites, as reflected by the over 25 different naming conventions among just 96 cookie control links we identified on websites (see Table LABEL:tab:cookie-control-name in the Appendix)—echoing prior work that documented similarly diverse cookie consent interfaces (Degeling et al., 2018; Utz et al., 2019). We found 38 distinct labels for cookies that users can opt out of, and 13 labels for cookies that cannot be opted out of (see Tables LABEL:tab:cookie-controllable and LABEL:tab:cookie-uncontrollable in the Appendix). Notably, banks’ cookie controls inconsistently applied 10 of these labels: for example, 6 banks categorized “functional” cookies as non-opt-out-able, while 22 others allowed users to opt out. This inconsistent labeling may contribute to user confusion, as prior work has shown that users often mistake “functional” cookies for “strictly necessary” ones (Habib et al., 2022). This lack of standardization in cookie labeling and opt-out availability undermines transparency and may mislead consumers about their data control.
Table 6. Disclosure about allowing third-party cookies, the availability of cookie controls, and third-party cookie practices (n=2073).

### v2

@@ 5.2. RQ2: Third-Party Sharing Statements
Regarding inconsistencies in marketing-related sharing statements across policies, we first discuss the GLBA statements (5.2.1). After presenting relevant statement categories we identified in other policies (5.2.2), we compare banks’ denying GLBA statements (5.2.3) and affirmative GLBA statements (5.2.4), each with opposite statements in other polices. Last, we point out common misleading language use for no-sharing (5.2.5).
@@ 5.2.1. GLBA third-party sharing statements
Of the seven sharing purposes prescribed in the GLBA model notice’s data-sharing table (see Figure 1(a)), we focused on the four relevant to third-party sharing: 1,542 banks (77% out of 2,002 banks that provided this row) stated they are sharing consumer personal information for marketing (marketing), 879 (43.9% out of 2,002) disclosed sharing with nonaffiliated financial companies for joint marketing (joint marketing), 329 (16.4% out of 1,070) disclosed sharing with affiliates for affiliates’ marketing (affiliate sharing), and 65 (3.2% out of 1,925) disclosed sharing with nonaffiliates for nonaffiliates’ marketing (nonaffiliate sharing). A few banks (36, 1.8% out of 2,002 banks with a GLBA notice) shared data for all four marketing purposes, 10.6% (214) for three of them; a third for two purposes (643, 32.1%) and another third shared for one of these purposes (743, 37.1%). Only 5.6% (112) indicated not sharing for any of the four purposes. These four GLBA-defined third-party sharing practices have nuanced differences, however, for consumers, they all describe scenarios in which their personal financial information is shared by their bank with other entities for marketing use.
Some banks include additional data-sharing disclosures in their GLBA notices’ OII box. Among the 718 banks that had a non-empty OII box, 332 included California-specific language. 300 of them provided statements further limiting data-sharing practices for California residents. The most common phrasing was: “ We will not share personal information with non-affiliates either for them to market to you or for joint marketing without your authorization. We will also limit our sharing of personal information about you with our affiliates to comply with all California privacy laws that apply to us.”—which suggests that without consent, the bank does not share for joint marketing or nonaffiliate marketing for Californians, though remaining vague on affiliate sharing. 20 (of 332) banks included minimal or opaque disclosures that mentioned CalFIPA but only stated “In accordance with California law, [bank] does not share personal information we collect except as permitted by law.” 55 (of 332) mentioned or linked to the bank’s CCPA policy (e.g., “Please visit our California Privacy Rights Act Policy for more information.”). These references at least acknowledge the existence of additional privacy policies, but still require consumers to navigate beyond the standardized data-sharing table in the GLBA notice, into a less prominent section of the notice, follow a link, and then read those additional policies to fully understand the bank’s data-sharing practices.
@@ 5.2.2. Third-party sharing statements in other privacy policy types
Thus, for the purpose of comparing across policies, we created statement categories based on whether a third-party sharing statement refers to personal information in general or to a specific data type, and whether it indicates sharing with external entities for marketing-related purposes. These categories include both statements allowing/describing third-party sharing (“yes”) and explicitly denying sharing (“no”). However, in contrast to GLBA disclosures that require explicit “no” statements, most banks’ policies are silent on sharing practices that the bank presumably does not engage in. [^5] See examples of the sharing statement types in our Zenodo repository: https://doi.org/10.5281/zenodo.17014519 We treat marketing and advertising as similar purposes, hereafter referred to as marketing.
@@
We assessed when banks indicated a “no” for each of four GLBA marketing-related sharing purposes (marketing, joint marketing, affiliate sharing, nonaffiliate sharing), how many banks simultaneously disclosed a “yes” to third-party sharing in related categories identified above. We found that many banks still disclosed varying amounts of related sharing in other policies despite negative indications in their GLBA notice (see Table 2).
@@
Inconsistencies regarding nonaffiliate sharing. 1,860 banks (92.9% of 2,002 banks with an analyzable GLBA notice) stated in their GLBA notice that they do not share with nonaffiliates. Yet 481 of them (26.6%) also had affirmative sharing statements in other policies under at least one category we identified. Among these 481 banks, about 30% stated sharing personal information for advertising purposes, more than two-thirds disclosed using third-party analytics cookies, and about 40% for third-party marketing cookies. About 20% disclosed sharing personal information for marketing in CCPA-related disclosures. Possibly less concerning but still an indication of nonaffiliate sharing, less than 20% stated sharing for analytics purposes. More than 10% disclosed sharing specific personal information types for marketing (e.g., “[w]e may also share your device’s physical location, combined with information about what advertisements you viewed and other information we collect, with our marketing partners […]”). A small percent, but still 17 banks in number, even disclosed selling data under CCPA (e.g., “In the preceding twelve (12) months, we have sold personal information”).
@@
These inconsistencies may reflect GLBA’s narrow scope of personal information as relating to financial information only, whereas other privacy policies often cover a broader range of practices, including data collected through a bank’s website and mobile apps. Under GLBA, a “no” to nonaffiliate sharing means that the bank does not share (financial) personal information with nonaffiliates for those nonaffiliates’ marketing purposes. However, this does not preclude sharing with nonaffiliates for other purposes like supporting the bank’s own marketing efforts. This discrepancy is concerning as it may mislead consumers. A consumer may conclude from the GLBA notice that a bank does not share their personal information with others, while in practice (with a 25.8% likelihood) the bank still shares lots of their personal information with third parties.
@@
Inconsistencies regarding affiliate sharing. A similar definitional gap exists for affiliate sharing, which under GLBA is confined to sharing personal information with affiliated companies (e.g., a bank’s affiliated investment or loans business) for their marketing. Among the 741 banks that stated they do not share for affiliate marketing in their GLBA notice, 23.2% of them (172 banks) still disclosed other data-sharing practices in other policies. More than two-thirds of these 172 banks disclosed that they use third-party analytics cookies, more than a third disclosed third-party marketing cookies, and about a quarter stated sharing for marketing.
@@
Inconsistencies regarding other sharing. We have seen above that when inconsistencies occur, most of them are between GLBA and third-party cookie disclosures. This pattern persists for “no” responses to GLBA-defined joint marketing and general marketing. Among the 1,123 banks that said they do not share for joint marketing in GLBA notice, 257 banks (22.9% of them) in fact stated that they do share personal information for marketing purposes in other privacy policies. For those with at least one inconsistency, about 70% banks stated allowing third-party analytics cookies, and a third for marketing cookies. About a quarter of them stated sharing personal information for marketing. A smaller group of 460 banks indicated “no” to sharing for the GLBA-defined marketing purpose, and 68 (14.8% of them) disclosed at least one sharing practice in other policies. Compared to other purposes, under joint marketing and marketing, we found smaller percentages of banks that had at least one inconsistency. Yet, any inconsistencies under these two purposes are striking, since they are about sharing for the bank’s own marketing purposes, a practice likely more common than nonaffiliate or affiliate sharing that is more narrowly defined under GLBA.
@@
Taken together, of the banks that indicated “no” to at least one GLBA-defined sharing purpose, 492 also disclosed at least one related data sharing practice as occurring in other policies. They account for 53.8% of the 915 banks that provided a GLBA notice and at least one additional policy. This suggests that even when banks claim not to share under GLBA categories, they are very likely to still engage in similar data-sharing practices for marketing. In sum, inconsistency between GLBA’s and other policies’ disclosures is widespread, which highlights coverage gaps and the lack of compatible clarity across privacy disclosures.
@@ 5.2.4. More restrictive third-party sharing than affirmative GLBA statements
Mkt 37 41 114 184 88 326
Joint-mkt 19 30 65 96 52 192
Aff-mkt 7 14 20 84 36 117
Nonaff-mkt 1 4 6 13 4 21
Some banks that indicated they share personal information under GLBA-defined categories simultaneously stated in their other privacy policies that they do not share information with third parties, most commonly in their CCPA-related sharing disclosures (see Table 3).
Among the 329 banks that disclosed affiliate sharing in their GLBA notice (16.4% of 2,002 banks with an analyzable GLBA notice), 117 banks (35.6%) had at least one statement denying data-sharing in their other policies. 71.8% these 117 banks stated that they do not share personal information, with or without reference to CCPA’s definition of sharing in CCPA policies or California-specific sections, and 30.8% of them stated that they don’t sell personal information in CCPA-related disclosures.
Regarding the GLBA’s nonaffiliate sharing, similarly high percentages of banks disclosed at least one denying data-sharing statement in their other policies, though fewer banks in number did so. Among the 65 banks (3.2% of 2,002 banks with an analyzable GLBA notice) that stated a “yes” to nonaffiliate sharing, 30.1% of them (21 banks) disclosed a denial of data-sharing of some kind. 62% of the 21 banks indicated that they do not share, and 19% do not sell, in CCPA-related disclosures.
Among banks that indicated a “yes” to GLBA’s marketing or joint-marketing purposes, smaller percentages (about 20% for each), though with larger counts, denied respective data-sharing elsewhere. For both, more than half of the ones with inconsistencies disclosed not sharing or selling in CCPA-related disclosures.
In total, 337 banks that disclosed sharing under GLBA indicated in other policies that they do not share/sell in CCPA-related sections. This pattern of discrepancy between GLBA and CCPA-related disclosures may indicate that banks specifically restrict their third-party sharing practices for their Californian customers while sharing their other customers’ data more freely.
@@ 5.2.5. Misleading “we don’t share” statements
Figure 4. Data-sharing statements and indication of whether consumers can limit the corresponding sharing practice in GLBA notices (n=2,002).
RQ2 summary. Banks’ affirmative sharing disclosures under the four GLBA marketing-related categories ranged widely, from 77% to just 3%. Comparing these with disclosures in other privacy policies, we noted two prevalent yet nuanced inconsistency types: (1) 492 banks (53.8% of those with both a GLBA and other policies) said “no” under GLBA while saying “yes” elsewhere for related sharing. This highlights that GLBA “no” statements alone are insufficient for understanding banks’ data practices, especially when online and mobile services are involved. (2) In contrast, 337 banks (36.8% of those with both a GLBA and other policies) disclosed sharing under GLBA and yet restricted such practices specifically for California residents. Additionally, although not contributing to the inconsistency types identified above, many “we don’t share” statements involved broad legal qualifiers like “except as permitted by law,” which complicates the interpretation of such no-sharing disclosures and may mislead consumers.
@@ 5.3. RQ3: Third-Party Sharing Opt-Outs
Table 4. Summary of opt-out controls provided (n=2,073).
@@ 4 Other combinations 4 4
Under the GLBA, banks that share for nonaffiliate marketing, affiliate marketing, or affiliate credit-related purposes must provide respective opt-outs. We found that almost all banks followed these requirements and provided required GLBA opt-outs (see Figure 4). For sharing purposes without required opt-outs, few banks that disclosed sharing provided opt-outs (joint marketing (6.0%), affiliate transactions (4.5%), bank’s own marketing (2.9%), and everyday business (0.1%)). This indicates that banks generally only offer opt-out controls when mandated by law. There is also a risk that banks may categorize their data sharing practices as “joint marketing” rather than sharing for affiliate or nonaffiliate marketing to sidestep GLBA opt-out obligations.
Furthermore, we found that banks’ GLBA-related opt-out options are often burdensome for consumers to exercise. Calling the bank was the most commonly mentioned opt-out method (307 of 367 banks providing GLBA-related opt-outs, 83.7%), and for many, the only method provided (114, 31.1% of 367). Additionally, 43 banks (11.7% of 367) required consumers to submit opt-out requests by mail, with no other methods offered. Only a third of banks (139, 37.9% of 367) offered a link to their webpage. Thus, while most banks technically comply with GLBA opt-out requirements, the opt-out methods are often unnecessarily cumbersome, potentially discouraging consumers from exercising their privacy rights.
@@ 5.3.3. Third-party cookie controls
300 banks’ privacy policies mentioned that consumers can opt out of their use of cookies, but through browsers, e.g., “You can set your browser to refuse Cookies.” 106 banks in policies referred consumers to cookie opt-outs offered by third parties (e.g., the Network Advertising Initiative), often for advertising cookies specifically. Only 17 banks mentioned using their cookie setting/banner as an opt-out choice in policies.
We examined cookie controls implemented/provided on websites. Among the 332 banks that indicated allowing third-party cookies for marketing, analytics, or both purposes in policies, only 43 (13.0%) of them implemented cookie controls (see Table 6). On the other hand, among the vast majority of banks (1,741) that did not disclose cookie practices in privacy policies, 53 (3.0% of 1,741) still implemented cookie controls. This suggests that they are engaging in cookie-related data sharing, which potentially involves third-party sharing, despite not explicitly disclosing it.
The 96 cookie control links we identified were referred to by 25 different names, echoing prior work that documented similarly diverse cookie consent interfaces (Degeling et al., 2018; Utz et al., 2019). We also found 38 distinct labels for cookies that users can opt out of, and 13 labels for cookies that users cannot. Banks’ cookie controls inconsistently applied 10 of these labels: for example, 6 banks categorized “functional” cookies as non-opt-out-able, while 22 banks allowed users to opt out. This inconsistent labeling may contribute to user confusion, as prior work has shown that users often mistake “functional” cookies for “strictly necessary” ones (Habib et al., 2022). Taken together, the lack of opt-out availability and standardization in cookie labeling undermines transparency and may mislead consumers about their choices.
Table 6. Disclosure about allowing third-party cookies, the availability of cookie controls, and third-party cookie practices (n=2,073).